# Multi-directional investigations on quiet time suprathermal ions measured by ASPEX-STEPS on-board Aditya L1

### v1

@@ 3.2 Estimation of spectral index
We present below (Figure 3) typical examples of ion spectra obtained from PS, IM, EP and NP sensors.
Figure 3: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for the PS (a), IM (b), EP (c) and NP (d) directions of AL1-ASPEX-STEPS. Spectral indices are calculated by fitting only the linear part in this log-log plots. The range of energy for the fitted linear part is mentioned in each subplot for PS, IM, EP and NP in red, blue, green and magenta colors respectively.
In figure 3, we show the typical Ion spectra for PS, IM, EP and NP sensors during 07 March, 2024 to 10 March 2024. As can be seen, the nature of the spectra is different at the lower energies. This is due to low-level discriminator (LLD) cut-off at the instrument level. This is due to low-level discriminator (LLD) threshold, which is set electronically to reduce the instrument background noise. The counts below this LLD threshold are considered to be zero and above this value, there is non-linear behaviour up to a certain energy as seen in Figure 3. We fit $J=AE^{-m}$ function to the linear portions of the spectra to calculate the spectral indices. The linear portion of the spectra (in log-log curve) for PS, IM, EP and NP sensors are found to be within the energy ranges of 0.36-1.32 MeV, 0.14-1.22 MeV, 0.39-1.33 MeV and 0.12-1.23 MeV respectively. The energy bins are calculated within these linear fit regions. Thereafter, spectral indices along with the standard errors (fitting errors) are calculated using these energy bins. In this analysis, it is to be noted that the counting uncertainties are also considered while estimating the spectral indices. The counting uncertainties in the flux values are shown as vertical error bars. Since the vertical error bars are very small, these are not clearly visible in log-log scale.
@@ 4 Results
In this section, the spectral indices are calculated for all the four ASPEX-STEPS sensors (PS, IM, EP, and NP) under three different scenarios. As the quiet intervals are not always identical for all the four sensors, in section 4.1, we find out common quiet intervals and derive the spectral indices for the four sensors. In section 4.2, we focus on individual sensors and consider all the quiet intervals for that sensor. Further, we also consider a period when the Aditya-L1 spacecraft was rotated for some time and therefore, the four STEPS sensors also rotated from their nominal positions during this interval. This scenario (section 4.3) is brought in to verify whether the spectral indices derived in the first two scenarios remain consistent in the third scenario. The spectral indices derived for these three scenarios are presented in the ensuing subsections.
4.1 Common quiet intervals for PS, IM, EP and NP sensors
@@
We present in Figure 4 the spectral fits and the corresponding indices calculated for a few common quiet intervals recorded by the PS, IM, EP and NP sensors. The duration of these intervals ranges from a day to a few days as can be seen from the top of each of the subplots. Each row in Figure 4 represents a particular sensor but for 5 different intervals. The values of spectral indices (standard fitting errors are also mentioned) for common quiet intervals are for all the sensors are found to be nearly similar, i.e. $\sim 2.0$, suggesting nearly identical spectral indices in all directions.
@@
Figure 4: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for the PS (a-e), IM (f-j), EP (k-o) and NP (p-t) sensors of AL1-ASPEX-STEPS.
4.2 Quiet intervals for individual sensors (PS, IM, EP and NP)
@@
In this section, we present (Figures 5-8) the spectral fits as well as indices calculated for the quiet intervals selected for each individual sensor. Similar to Figure 4, here also the quiet intervals are mentioned at the top of each sub-plots. The spectral indices calculated for quiet intervals of each sensor also show similar values of $\sim 2.0$. Therefore, regardless of the intervals and sensors, the spectral indices are found to be close to 2.
@@
Figure 5: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for the Parker Spiral (PS) sensor of AL1-ASPEX-STEPS.
@@
Figure 6: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for Intermediate (IM) sensor of AL1-ASPEX-STEPS.
@@
Figure 7: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for Earth pointing (EP) sensor of AL1-ASPEX-STEPS.
@@
Figure 8: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for North pointing (NP) sensor of AL1-ASPEX-STEPS.
4.3 Quiet intervals when AL1 was rotated (during November - 2024)
@@
During the quiet interval of 25 November 2024, Aditya L1 (AL1) spacecraft underwent a rotation for a few hours (06:30 UTC to 14:10 UTC). In this rotation period, PS, IM, EP and NP detectors also rotated from their nominal positions. Through Figure 9, we illustrate the orientations of the PS, IM, NP, and EP sensors in their nominal and rotated configurations, represented in the Geocentric Solar Ecliptic (GSE) coordinate system. Panels (a–h) show the nominal configuration, while panels (i–p) display the rotated configuration. In all panels, the X-axis represents the direction cosine corresponding to the $X_{GSE}$ component for each of the four sensors. Panels (a–d) and (i–l) present the direction cosines of the $Y_{GSE}$ component along the Y-axis, whereas panels (e–h) and (m–p) show the direction cosines of the $Z_{GSE}$ component along the Y-axis. Each panel includes a blue and a yellow dot denoting the Earth and the Sun, respectively. The red and green arrows indicate the orientations of the PS, IM, NP, and EP sensors projected in the X–Y and X–Z planes of the GSE coordinate system, respectively. The length of each arrow is proportional to the corresponding direction cosine component, thus accurately reflecting the detector orientation.
@@
Figure 9: Direction cosines of PS, IM, NP and EP sensors in nominal configuration (panel a-h) and in rotated configuration (panel i-p) of AL1-ASPEX-STEPS. In subplots (a-p) blue dot and yellow dot represents the Earth and the Sun respectively. Red and green arrows represent the directions of PS, IM, NP and EP sensors in XY and XZ plane of GSE coordinate system respectively. The length of an arrow is decided by the corresponding values of direction cosines.
We estimated the spectral indices for the PS, IM and NP sensors during this quiet interval when the spacecraft was rotated (shown in Figure 10). For the EP sensor, the quiet time criteria did not get satisfied and hence spectral indices are not calculated for the EP sensor during this period. Interestingly, spectral indices calculated during this interval also show the similar values $\sim 2.0$ regardless of the sensors and directions.
@@
Figure 10: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for PS, IM and NP sensors of AL1-ASPEX-STEPS during rotated period.
In summary, the Figures 4-10 reveal that spectral indices are close to 2 regardless of the direction and sensor during quiet intervals.
@@ 5 Discussions
In this study, we have calculated spectral indices for quiet time suprathermal ions in IP medium using measurements from the four sensors of STEPS instrument of ASPEX onboard Aditya L1 spacecraft for different quiet intervals as well as during nominal and rotated configurations of the spacecraft. The value of spectral indices obtained from PS, IM, EP and NP sensors are found to be $\sim 2.0$ that suggests isotropic distribution of quiet time suprathermal ions. A few earlier studies (e.g. M. Dayeh et al. 2009, 2017; B. Dalal et al. 2022) have shown that the spectral indices vary over a wide range even for quiet periods. These studies were based on the longer intervals of data obtained from instruments onboard spinning spacecraft such as Wind and Advanced Composition Explorer (ACE). Our analysis is done for shorter intervals of measurements in a directionally resolved manner. In order to assess whether this apparent isotropy reflects a genuine physical property of the ST ions distribution or is influenced by the relative motion between the spacecraft and the solar wind frame, a correction was applied using the Compton-Getting effect (F. Ipavich, 1974). The Compton-Getting effect accounts for the alteration in the observed energy spectra of particles due to the Doppler-like shift arising from the motion of the observer with respect to the solar wind rest frame. This correction enables a more accurate inference of the intrinsic particle distribution by eliminating systematic effects related to the observer’s velocity. After considering the Compton-Getting effect, we obtain the corrected or effective spectral index, which better represents the true properties of the particles.
@@
Spectral indices obtained by applying the correction due to the Compton- Getting effect do not exhibit the significant changes up to the first order and remains almost the same $\sim 2.0$. The reason behind the negligible changes in the spectral indices after applying Compton- Getting effect is that the energies (speeds) of the suprathermal ions are very high in comparison to relative speed $\left(w\right)$ between solar wind and spacecraft which makes the term $\left(\frac{w}{v}\right)$ negligible in equation 2. In our analysis, the angle $\theta$ between the solar wind velocity and the PS, IM, EP and NP direction is $52^{\circ}$, $30^{\circ}$, $156^{\circ}$, and $90^{\circ}$ respectively (for more details, refer S. K. Goyal et al. 2025). There will be no Compton-Getting correction in the North Pointing (NP) direction as $\theta=90^{\circ}$ which makes $\cos\theta$ term zero in equation 2. For all other angles, the Compton-Getting corrections are also minimal. Thus, negligible changes are seen in the values of spectral indices for PS, IM and EP directions after applying the Compton- Getting corrections. This shows that the isotropy is a genuine physical property of the suprathermal ions distribution and this is not influenced by the relative motion between the spacecraft and the solar wind frame. This investigation validates the assumptions of isotropic distribution of quiet time suprathermal ions made by L. Fisk & G. Gloeckler (2008) in solving Parker transport equation (PTE) for acceleration of suprathermal ions in the solar wind.
@@
Another important insight that comes out of this study is that the spectral indices remain nearly identical $\left(\sim 2\right)$ for quiet intervals if we consider a time scale of a few days. This is in contrast to previous studies that reported significant variability in spectral indices when large intervals were considered. For instance, M. Dayeh et al. (2009, 2017) calculated spectral indices using a one-year average for eighteen years (1998-2015) and observed a wide range of values of spectral indices for different elements. Similarly, B. Dalal et al. (2022) analyzed spectral indices using a two-year average across different phases of the solar cycle—such as solar maximum, declining phase, solar minimum, and rising phase—and also reported substantial variations in the spectral index values for different elements. This significant variability in spectral index values indicates that even during the quiet period, different processes dominate in different phases of solar cycles. M. Dayeh et al. (2017) suggests that while impulsive and shock acceleration events dominate during enhanced solar activity periods, stream/corotating interaction regions (SIR/CIR) and heated solar wind processes dominate during low solar activity periods. These authors also suggest that the large variability of spectral indices above $100keV$ indicate towards the dominant role of these large, transient processes. Earlier studies by Y. Chen et al. (2015) and R. Hajra & J. V. Sunny (2022) have shown that SIR/CIRs are more prevalent during the declining and minimum phases of the solar cycle. B. Dalal et al. (2023) analyzed the spectral indices of suprathermal ions associated with multiple SIR/CIRs and found typical values of spectral indices in the range of 3-4, indicating significantly softer spectra. In contrast, our study, conducted during a solar maximum period, reports spectral indices closer to 2 in all the four directions. This suggests that during solar minimum, the quiet time suprathermal ion pool is primarily influenced by ions accelerated by CIRs/SIRs, resulting in softer spectra while during solar maximum, the suprathermal ion population during quiet periods is possibly dominated by large, impulsive and transient events like impulsive and gradual solar energetic particle events (SEPs), leading to harder spectral indices. In this context, the time scale of a few days is also important. We believe that the influence of the large transient processes is minimized by this choice leading to the nearly identical spectral indices recorded by all the four sensors of AL1-ASPEX-STEPS. As 2024 is a high solar activity year and AL1-ASPEX-STEPS will continue to make measurements, it will be interesting to see whether this nearly identical spectral index continue to hold during low solar activity period as well and if the magnitude of spectral index changes. Further, since ASPEX-STEPS does not discriminate between species and primarily measures protons and alpha particles (being the most abundant), the $m/q$ dependences of the acceleration processes are also averaged out. This suggests that it is difficult to comment on the individual processes that fix the spectral indices close to 2 at shorter time scales based on the present set of measurements. Therefore, in order to investigate the dominant processes influencing the isotropic distribution of quiet time suprathermal ions for the same periods as those of ASPEX-STEPS measurements, elemental abundance ratios were estimated using data from the Ultra-Low Energy Isotope Spectrometer (ULEIS) onboard the Advanced Composition Explorer (ACE) spacecraft. This is done following the same methodology suggested in the work of M. Dayeh et al. (2009, 2017).
@@
Insights drawn from Table 1 reveal notable signatures of impulsive solar energetic particle (ISEP) contributions during several intervals that are otherwise considered quiet. Specifically, substantial enhancements in the ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$ ratio commonly associated with ISEPs are observed during periods such as 04–15 March, 05–17 July, 02–04 October, 16–24 October, 14–21 November, and 23–30 November 2024. These elevated ${}^{3}\mathrm{He}/{}^{4}\mathrm{He}$ values, reaching or exceeding 0.1, suggest that approximately 25% of the selected quiet intervals are influenced by residual impulsive SEPs from previous events. In addition, the $\mathrm{Fe}/\mathrm{O}$ ratio exhibits considerable variability throughout the year. Notably, elevated $\mathrm{Fe}/\mathrm{O}$ values greater than 2, as seen on 28 January and during 13–15 August 2024, further support the presence of ISEP-like signatures. In contrast, approximately 45% of the quiet periods display $\mathrm{Fe}/\mathrm{O}$ ratios near unity (Fe/O $\sim 1$), indicating a substantial contribution from impulsive solar energetic particles (ISEPs) events, which are characterized by enhanced heavy ion abundances. The $\mathrm{C}/\mathrm{O}$ ratio adds further insight into the compositional makeup of the suprathermal ion population. Approximately 20% of the quiet intervals show $\mathrm{C}/\mathrm{O}$ values near 0.32, indicative of impulsive SEP contributions, while around 50% of the intervals display values near 0.42, which are characteristic of gradual SEP populations. Together, these findings underscore the complexity of particle composition even during nominally quiet times and highlight the persistent influence of impulsive events in shaping suprathermal ion populations.

### v2

@@ 3.2 Estimation of spectral index
Figure 3: Plots of quiet time differential directional flux versus energy with spectral index and coefficient of determination ($R^{2}$) values mentioned in each subplot for the PS (a), IM (b), EP (c) and NP (d) directions of AL1-ASPEX-STEPS. Spectral indices are calculated by fitting only the linear part in this log-log plots. The range of energy for the fitted linear part is mentioned in each subplot for PS, IM, EP and NP in red, blue, green and magenta colors respectively.
In figure 3, we show the typical Ion spectra for PS, IM, EP and NP detectors during 07 March, 2024 to 10 March 2024. As can be seen, the nature of the spectra is different at the lower energies. This is due to low-level discriminator (LLD) threshold, which is set electronically to reduce the instrument background noise. The counts below this LLD threshold are considered to be zero and above this value, there is non-linear behaviour up to a certain energy as seen in Figure 3. Towards the low-energy end of the spectrum, the flux distribution exhibits an initial enhancement just above the threshold, followed by a fall before settling into the real physical trend. This initial rise arises due to the response of the solid-state detectors and electronics employing Schmitt trigger based Lower-Level Discriminator (LLD) circuit in the instrument. The LLD sets a minimum amplitude threshold on the detector signals to suppress electronic noise and very low-energy background counts, thereby ensuring reliable measurements of the real signals. However, due to the hysteresis in the values of the LLD cut off, spurious counts from incomplete charge collection and residual electronic noise still pass, leading to an apparent rise in flux values. Different values of low energy enhancements as seen in different detectors are further dependent on the detector types such as the PS and EP detectors, both of which have a dead layer of about $0.8~\mu\mathrm{m}$, display similar low-energy distortions, while the IM and NP detectors, with thinner dead layers of about $0.2~\mu\mathrm{m}$, show a comparable response pattern distinct from the PS and EP detectors. This effect has been further verified during in-orbit payload verification phase, by setting different LLD values for the detectors. At present, the final set LLD values are different for different detectors based on the electronics noise covering the possible lowest energy range. Thus, the apparent enhancement and fall observed at the lowest energies are instrument induced rather than of physical origin. Beyond this region as the energy increases, the signal becomes clean and influence of such artefacts diminishes, which subsequently follows the expected power-law trend. We fit $J=AE^{-m}$ function to the linear portions of the spectra to calculate the spectral indices. To determine the linear portion of the spectra, we evaluated the coefficient of determination ($R^{2}$) by systematically varying the lower bound of the fitting interval, starting from the point where the spectra visually exhibit a linear trend. The lower energy bound corresponding to the highest ($R^{2}$) value, obtained after fitting a power-law function, was adopted as the starting energy of the linear portion. The upper energy limits of linear portion of the spectra for the PS, IM, EP, and NP detectors were taken to be 1.32 MeV/n, 1.22 MeV/n, 1.33 MeV/n and 1.23 MeV/n respectively. Thus, the linear portion of the spectra (in log-log space) for PS, IM, EP and NP detectors are found to be within the energy ranges of 0.36-1.32 MeV/n, 0.14-1.22 MeV/n, 0.39-1.33 MeV/n and 0.12-1.23 MeV/n respectively. The energy bins are calculated within these linear fit regions. Thereafter, spectral indices along with the standard errors (fitting errors) are calculated using these energy bins. In this analysis, it is to be noted that the counting uncertainties are also considered while estimating the spectral indices. The counting uncertainties in the flux values are shown as vertical error bars. Since the vertical error bars are very small, these are not clearly visible in log-log scale.
@@ 4 Results
In this section, the spectral indices are calculated for all the four ASPEX-STEPS detectors (PS, IM, EP, and NP) under three different scenarios. As the quiet time are not always identical for all the four detectors, in section 4.1, we identify common quiet time and derive the spectral indices for the four detectors. In section 4.2, we focus on individual detectors and consider all the quiet time for that detector. Further, we also consider a period when the Aditya-L1 spacecraft was rotated for some time and therefore, the four STEPS detectors also rotated from their nominal positions during this interval. This scenario (section 4.3) is brought in to verify whether the spectral indices derived in the first two scenarios remain consistent in the third scenario. The spectral indices derived for these three scenarios are presented in the ensuing subsections.
4.1 Common quiet time for PS, IM, EP and NP detectors
@@
We present in Figure 4 the spectral fits and the corresponding indices calculated for a few common quiet time recorded by the PS, IM, EP and NP detectors. The duration of these intervals ranges from a day to a few days as can be seen from the top of each of the plots. Each row in Figure 4 represents a particular detector but for 5 different intervals. The values of spectral indices (with standard fitting errors) for the common quiet-time interval across all detectors are found to be nearly similar, within the range of $1.99\pm 0.06$ to $2.15\pm 0.02$ (average value $2.07\pm 0.04$), suggesting nearly identical spectral indices in all directions.
@@
Figure 4: Plots of quiet-time differential directional flux versus energy with spectral index values shown in each subplot for the PS (a–e), IM (f–j), EP (k–o), and NP (p–t) detectors of AL1-ASPEX-STEPS.
4.2 Quiet time for individual detectors (PS, IM, EP and NP)
@@
In this section, we present (Figures 5-8) the spectral fits as well as indices calculated for the quiet time selected for each individual detector. Similar to Figure 4, the quiet time are listed at the top of each plots. The spectral indices calculated for quiet time of each detector also show similar values within the range of $1.84\pm 0.09$ to $2.16\pm 0.02$ (average value $2.0\pm 0.06$). Therefore, regardless of the intervals and detectors, the spectral index values are found to be similar and close to $\sim-2.0$.
@@
Figure 5: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for the Parker Spiral (PS) detector of AL1-ASPEX-STEPS.
@@
Figure 6: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for Intermediate (IM) detector of AL1-ASPEX-STEPS.
@@
Figure 7: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for Earth pointing (EP) detector of AL1-ASPEX-STEPS.
@@
Figure 8: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for North pointing (NP) detector of AL1-ASPEX-STEPS.
4.3 Quiet time when AL1 was rotated (during November - 2024)
@@
During the quiet interval of 25 November 2024, Aditya L1 (AL1) spacecraft underwent a rotation for a few hours (06:30 UTC to 14:10 UTC). In this rotation period, PS, IM, EP and NP detectors also rotated from their nominal positions. Through Figure 9, we illustrate the orientations of the PS, IM, NP, and EP detectors in their nominal and rotated configurations, represented in the Geocentric Solar Ecliptic (GSE) coordinate system. Panels (a–h) show the nominal configuration, while panels (i–p) display the rotated configuration. In all panels, the X-axis represents the direction cosine corresponding to the $X_{GSE}$ component for each of the four detectors. Panels (a–d) and (i–l) present the direction cosines of the $Y_{GSE}$ component along the Y-axis, whereas panels (e–h) and (m–p) show the direction cosines of the $Z_{GSE}$ component along the Y-axis. Each panel includes a blue and a yellow dot denoting the Earth and the Sun, respectively. The red and green arrows indicate the orientations of the PS, IM, NP, and EP detectors projected in the X–Y and X–Z planes of the GSE coordinate system, respectively. The length of each arrow is proportional to the corresponding direction cosine component, thus accurately reflecting the detector orientation. We estimated the spectral indices for the PS, IM and NP detectors during this quiet interval when the spacecraft was rotated (shown in Figure 10). For the EP detector, the quiet time criteria did not get satisfied and hence spectral indices are not calculated for the EP detector during this period. Interestingly, spectral indices calculated during this interval also show the similar values within the range of $1.99\pm 0.04$ to $2.12\pm 0.03$ (average value $2.05\pm 0.04$) regardless of the detectors and directions. In summary, the Figures 4-10 reveal that spectral indices are almost similar ($\sim-2.0$) regardless of the direction and detector during quiet time.
@@
Figure 9: Direction cosines of PS, IM, NP and EP detectors in nominal configuration (panel a-h) and in rotated configuration (panel i-p) of AL1-ASPEX-STEPS. In plots (a-p) blue dot and yellow dot represents the Earth and the Sun respectively. Red and green arrows represent the directions of PS, IM, NP and EP detectors in XY and XZ plane of GSE coordinate system respectively. The length of an arrow is decided by the corresponding values of direction cosines.
@@
Figure 10: Plots of quiet time differential directional flux versus energy with spectral indices value mentioned in each subplot for PS, IM and NP detectors of AL1-ASPEX-STEPS during rotated period.
@@ 5 Discussions
In this study, we have calculated spectral indices for quiet time suprathermal ions in IP medium using measurements from four detectors on STEPS instrument of ASPEX onboard Aditya L1 spacecraft during nominal and rotated configurations of the spacecraft. The value of spectral indices obtained from PS, IM, EP and NP detectors are found to be within the range of $1.84\pm 0.09$ to $2.15\pm 0.02$ (average value $1.99\pm 0.06$) that suggests isotropic distribution of quiet time suprathermal ions. A few earlier studies (e.g. M. Dayeh et al. 2009, 2017; B. Dalal et al. 2022) have shown that the spectral indices vary over a wide range even for quiet periods. These studies were based on the longer intervals of data obtained from instruments onboard spinning spacecraft such as Wind and Advanced Composition Explorer (ACE). Our analysis is done for shorter intervals of measurements in a directionally resolved manner. In order to assess whether this apparent isotropy reflects a genuine physical property of the ST ions distribution or is influenced by the relative motion between the spacecraft and the solar wind frame, a correction was applied using the Compton-Getting effect (F. Ipavich, 1974). The Compton-Getting effect accounts for the alteration in the observed energy spectra of particles due to the Doppler-like shift arising from the motion of the observer with respect to the solar wind rest frame. This correction enables a more accurate inference of the intrinsic particle distribution by eliminating systematic effects related to the observer’s velocity. After considering the Compton-Getting effect, we obtain the corrected or effective spectral index, which better represents the true properties of the particles.
@@
Spectral indices obtained by applying the correction due to the Compton- Getting effect do not exhibit significant changes up to the first order and remains almost the same $\sim-2.0$. The reason behind the negligible changes in the spectral indices after applying Compton- Getting effect is that the energies (speeds) of the suprathermal ions are very high in comparison to relative speed $\left(w\right)$ between solar wind and spacecraft which makes the term $\left(\frac{w}{v}\right)$ negligible in equation 2. In our analysis, the angle $\theta$ between the solar wind velocity and the PS, IM, EP and NP direction is $52^{\circ}$, $30^{\circ}$, $156^{\circ}$, and $90^{\circ}$ respectively (for more details, refer S. K. Goyal et al. 2025). There will be no Compton-Getting correction in the North Pointing (NP) direction as $\theta=90^{\circ}$ which makes $\cos\theta$ term zero in equation 2. For all other angles, the Compton-Getting corrections are also minimal. Thus, negligible changes are seen in the values of spectral indices for PS, IM and EP directions after applying the Compton- Getting corrections. This shows that the isotropy is a genuine physical property of the suprathermal ions distribution and this is not influenced by the relative motion between the spacecraft and the solar wind frame. This investigation validates the assumptions of isotropic distribution of quiet time suprathermal ions made by L. Fisk & G. Gloeckler (2008) in solving Parker transport equation (PTE) for acceleration of suprathermal ions in the solar wind.
@@
Another important insight that comes out of this study is that the spectral indices remain nearly identical $\left(\sim-2\right)$ for quiet time if we consider a time scale of a few days. This is in contrast to previous studies that reported significant variability in spectral indices when large intervals were considered. For instance, M. Dayeh et al. (2009, 2017) calculated spectral indices using a one-year average for eighteen years (1998–2015) and observed a wide range of values of spectral indices for different elements. Similarly, B. Dalal et al. (2022) analyzed spectral indices using a two-year average across different phases of the solar cycle such as solar maximum, declining phase, solar minimum, and rising phase—and also reported substantial variations in the spectral index values for different elements. Such variability arises because the degree of acceleration experienced by suprathermal populations can vary depending on species, since it is dependent on $m/q$ (B. Dalal et al., 2022). Quiet-time suprathermal ion pools may also contain remnants of earlier transient events (e.g., SEPs), leading to variability in spectral indices as the abundance of different species at a given place and time may depend on the type of SEP event and their $m/q$ dependent transport along interplanetary magnetic field lines. Moreover, the relative dominance of transient solar and interplanetary drivers such as flares, ICMEs, and SIRs that supply suprathermal ions in the interplanetary medium changes with solar phase. This can cause shifts in the dominant acceleration mechanism, producing variability in the spectral indices of quiet-time suprathermals when long intervals are considered. An additional factor is the definition of threshold levels: earlier works (e.g., M. Dayeh et al. (2017); B. Dalal et al. (2022)) calculated thresholds based on one- or two-year averages, but such values themselves can vary across solar phases and cycles. This significant variability in spectral index values therefore indicates that even during the quiet period, different processes dominate in different phases of solar cycles. M. Dayeh et al. (2017) suggest that while impulsive and shock acceleration events dominate during enhanced solar activity periods, stream/corotating interaction regions (SIR/CIR) and heated solar wind processes dominate during low solar activity periods. These authors also suggest that the large variability of spectral indices above $100$ keV/n indicates the dominant role of such large, transient processes. Earlier studies by J. Gosling (1996), Y. Chen et al. (2015) and R. Hajra & J. V. Sunny (2022) have shown that SIR/CIRs are more prevalent during the declining and minimum phases of the solar cycle. B. Dalal et al. (2023) analyzed the spectral indices of suprathermal ions associated with multiple SIR/CIRs and found typical values of spectral indices in the range of 3-3 to 4–4, indicating significantly softer spectra. In contrast, our study, conducted during a solar maximum period, reports spectral indices closer to 2-2 in all four directions. This suggests that during solar minimum, the quiet-time suprathermal ion pool is primarily influenced by ions accelerated by CIRs/SIRs (J. Gosling, 1996), resulting in softer spectra (B. Dalal et al., 2023), while during solar maximum, the suprathermal ion population during quiet periods is dominated by large, impulsive and transient events such as solar energetic particle events (SEPs) (M. Dayeh et al., 2009, 2017), leading to harder spectral indices. In this context, the time scale of a few days is also important. By focusing on short time averaging, fixing a year within a solar cycle, and using species-integrated fluxes, we minimize the variable factors mentioned above, leading to the nearly identical spectral indices recorded by all four detectors of AL1-ASPEX-STEPS. As 2024 is a high solar activity year and AL1-ASPEX-STEPS will continue to make measurements, it will be interesting to see whether this nearly identical spectral index continues to hold during low solar activity periods as well, and if the magnitude of the index changes. Further, since ASPEX-STEPS does not discriminate between species and primarily measures protons and alpha particles (being the most abundant), the $m/q$ dependence of the acceleration processes is effectively averaged out. This makes it difficult to comment on the individual processes that fix the spectral indices close to 2-2 at shorter time scales based on the present set of measurements. Therefore, in order to investigate the dominant processes influencing the isotropic distribution of quiet-time suprathermal ions for the same periods as those of ASPEX-STEPS measurements, elemental abundance ratios were estimated using data from the Ultra-Low Energy Isotope Spectrometer (ULEIS) onboard the Advanced Composition Explorer (ACE) spacecraft. This is done following the same methodology suggested in the work of M. Dayeh et al. (2009, 2017).
@@
Table 1 reveals characteristic signatures of impulsive solar energetic particle (ISEP) contributions during several intervals that are otherwise considered quiet. Specifically, substantial enhancements in the $^{3}\mathrm{He}/{}^{4}\mathrm{He}$ ratio are observed during many quiet intervals that are commonly associated with ISEPs. Some of these intervals are 04–15 March, 05–17 July, 02–04 October, 16–24 October, 14–21 November, and 23–30 November 2024. The elevated $^{3}\mathrm{He}/{}^{4}\mathrm{He}$ values during these intervals, reaching or exceeding 0.1, suggest that approximately 25% of the selected quiet time are influenced by residual impulsive SEPs from previous events. In addition, the $\mathrm{Fe}/\mathrm{O}$ ratio exhibits considerable variability throughout the year. Notably, elevated $\mathrm{Fe}/\mathrm{O}$ values greater than 2, as seen on 28 January and during 13–15 August 2024, provides strong indication on the presence of ISEP remnants in the suprathermal ion pool. It is seen that approximately 45% of the quiet periods display $\mathrm{Fe}/\mathrm{O}$ ratios near unity (Fe/O $\sim 1$), indicating the contribution of ISEP events, which are characterized by enhanced heavy ion abundances. The $\mathrm{C}/\mathrm{O}$ ratio adds further insight into the compositional makeup of the suprathermal ion population. Approximately 20% of the quiet time show $\mathrm{C}/\mathrm{O}$ values near 0.32, indicative of impulsive SEP contributions, while around 50% of the intervals display values near 0.42, which are characteristic of gradual SEP populations. Together, these findings underscore the complexity of particle composition even during nominally quiet times and highlight the persistent influence of impulsive events in determining the compositional makeup of the suprathermal ion populations.